a gaseous compound contain 33% silicon and 67% fluorine by mass, at 35 degree. 0.210L of the compound exert a pressure of 0.2atm. if the mass of the 0.210l of the compound was 2.38g,

1) calculate the empirical formula of the compound

2) calculate the relative molecular mass of the compound

3) calculate the molecular formula of the compound

1
Expert's answer
2021-01-11T03:49:38-0500


nSi = 33.0g Si x "\frac{1mol Si}{28.09gSi}" = 1.17 mol Si ÷ 1.17 = 1 mol Si

nF = 67.0g F x "\frac{1molF}{19.00gF}" = 3.53 mol F ÷ 1.17 = 3 mol F

empirical formula is "\implies" SiF3


relative molecular mass of the compound SiF3

= 28.09 + 3(19) = 28.09 + 57 = 85.09 = 85.1 g


with volume and mass we can find gas density to solve for the molar mass

M = "\frac{dRT}{P}"

density of gas = 2.38 g ÷ 0.210 L = 11.33 g/L

M = "\frac{dRT}{P}" = "\frac{(11.33g/L)(0.0821)(308K)}{0.2atm}" = 1432.49 g /mol


empirical formula SiF3 = mm =85.1 g

"\frac{M}{mm}" = "\frac{1432.49}{85.1}" 16


molecular formula is Si16F48

molecular mass = 1361.44 g